OpenStudy (belleuuuuuuuuu):

2. Consider 2 1/2 divided by 1/4. (a) Write a real-world problem for the division. (b) Create and explain a model for the division. (c) Find the quotient for the real-world problem in part (a). Show your work or explain your reasoning. please help!!!

OpenStudy (studygurl14):

An easy real-world application of fractions is with baking. Say you're making breakfast for your whole family, who LOVES chocolate chip pancakes. But you only have 2 1/2 cups of chocolate chips, and you need 1/4 cup to make 1 batch of pancakes. How many batches can you make?

OpenStudy (studygurl14):

@belleuuuuuuuuu

OpenStudy (belleuuuuuuuuu):

5/8?

OpenStudy (belleuuuuuuuuu):

@StudyGurl14

OpenStudy (studygurl14):

First, convert 2 1/2 cups to a mixed fraction.

OpenStudy (belleuuuuuuuuu):

i got 5/2

OpenStudy (belleuuuuuuuuu):

@StudyGurl14

OpenStudy (studygurl14):

perfect. Now divide that by 1/4. Remember that when dividing by a fraction, you multiply it by its reciprocal

OpenStudy (belleuuuuuuuuu):

i divided 5/2 by 1/4 and i got 10/1

OpenStudy (studygurl14):

perfect. so the answer is 10

OpenStudy (belleuuuuuuuuu):

i know but my teacher said i have to Write a real-world problem for the division,Create and explain a model for the division and Show my work or explain your reasoning

OpenStudy (studygurl14):

Okay, remember my chocolate-chip pancake example?

OpenStudy (belleuuuuuuuuu):

yea

OpenStudy (studygurl14):

Okay, so this is a real world problem: You need to make chocolate-chip pancakes for your entire extended family. But wait, you only have 2 1/2 cups of chocolate chips. You need 1/4 cup of chocolate chips per batch of pancakes. How many batches can you make? \(\Large 2\frac{1}{2}\div\frac{1}{4}\) Step 1: \(\Large\frac{5}{2}\div\frac{1}{4}\) Step 2: \(\Large\frac{5}{2}\times\frac{4}{1}\) Step 3: \(\Large\frac{20}{2}\) Step 4: \(=10\) Luckily, you figured out you can make 10 batches of pancakes, which should be more than enough to feed your hungry family!

OpenStudy (belleuuuuuuuuu):

ohhh thank u so much

OpenStudy (studygurl14):

You're welcome. :)